OpenStudy (anonymous):

HELP ASAP!!!!!! What is the solution of the equation? 3^5 sqrt (x+2)^3 + 3 = 27???

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm Hello! }\) is this your problem \(\large\color{slate}{ 3\sqrt[5]{x+2)^2+3}=27 }\) ?

OpenStudy (anonymous):

Yep that's it.

OpenStudy (solomonzelman):

~ divide both sides by 3 ~ raise both sides to the fifth power

OpenStudy (anonymous):

Ok one sec

OpenStudy (solomonzelman):

take as much time as you need:)

OpenStudy (anonymous):

so once I do both of those, then would it look like this? sqrt(x+2)^3 + 3 = 27?

OpenStudy (solomonzelman):

yes, when you divide both sides by 3, you get: \(\large\color{slate}{ \sqrt[5]{x+2)^2+3}=9 }\)

OpenStudy (solomonzelman):

raise both sides to the 5th power.

OpenStudy (solomonzelman):

\(\large\color{slate}{ (x+2)^2+3=9^5 }\)

OpenStudy (solomonzelman):

can you finish from here?

OpenStudy (solomonzelman):

I got to go, actually sorry.... tag someone else for help

OpenStudy (anonymous):

Yes thank you SO much I really appreciate it I've been stuck on this for hours

OpenStudy (solomonzelman):

sure..